OpenStudy (pottersheep):

Lowest common denominator help pleaseee 3 / (x-2) - 2x/(x-2)^2

OpenStudy (anonymous):

lowest common denominator is \[(x-2)^2\]

OpenStudy (anonymous):

so if you want \[\frac{3}{x-2}-\frac{2x}{(x-2)^2}\] write \[\frac{3(x-2)-2x}{(x-2)^2}\] and then simplify the numerator

OpenStudy (pottersheep):

so you just switched the signs around for (x-2) ^2

OpenStudy (anonymous):

no i didn't switch the signs. if i want the lowest common denominator for 3 and 9 i take 9 if i want it for 5 and 25 use 25 and for 7 and 49 it is 49 in other words the lcm of x and x^2 is x^2 and likewise the lcm of (x-2) and (x-2)^2 is (x-2)^2

OpenStudy (pottersheep):

I wrote the question wrong, sorry. it was supposed to be 3 / (x-2) - 2x/(2-x)^2